OpenStudy (anonymous):

Solve the following equation: 4x2 + 13x = –3.

OpenStudy (anonymous):

solve it as 0

OpenStudy (anonymous):

4x^2 +13x +3 =0 4x^2 +12x+x+3 =0

OpenStudy (anonymous):

4x(x+3) +1(x+3) =0 (x+3)(4x+1) =0 x=-3 or x=-1/4

OpenStudy (anonymous):

got the same but ur quicker :)